# Turbocharging constraints on dark matter substructure through a synthesis of strong lensing flux ratios and extended lensed arcs

### 05 Mar 2024 v1

@@ 1 Introduction
In search of new physics, observational probes of dark matter structure from dwarf galaxies (Kim et al., 2018; Correa, 2021; Nadler et al., 2021; Bechtol et al., 2022; Dekker et al., 2022; Slone et al., 2023; Nadler et al., 2024), stellar streams (Bovy et al., 2017; Banik et al., 2018; Bonaca et al., 2019; Banik et al., 2021), and gravitational lensing (Dalal & Kochanek, 2002; Vegetti et al., 2014; Nierenberg et al., 2014; Hezaveh et al., 2016; Birrer et al., 2017; Nierenberg et al., 2017; Vegetti et al., 2018; Gilman et al., 2020, 2021; He et al., 2022; Sengül et al., 2022; Wagner-Carena et al., 2023; Dike et al., 2023; Powell et al., 2023; Dhanasingham et al., 2023; Keeley et al., 2023; Gilman et al., 2023; Mondino et al., 2023; Nightingale et al., 2024) characterize the properties of substructure on sub-galactic scales. If the dark matter has no coupling to the standard model besides gravity, these cosmic probes of dark matter constitute the only experiments with which to investigate its properties (see the reviews by Drlica-Wagner et al., 2022; Bechtol et al., 2022).
@@
We perform lensing calculations using the open-source software package lenstronomy [^1] https://github.com/lenstronomy/lenstronomy (Birrer & Amara, 2018; Birrer et al., 2021). We generate the populations of dark matter substructure using the open-source software pyHalo [^2] https://github.com/dangilman/pyHalo (Gilman et al., 2020). To perform the forward modeling calculations that interface between lenstronomy and pyHalo, we use the new open-source software samana [^3] https://github.com/dangilman/samana.
@@ 4.1 Modeling of dark matter subhalos and line-of-sight halos
where $\rho_{s}$ is a characteristic central density, $r_{s}$ is the scale radius, and $r_{t}$ implements a tidal radius. For field halos, we set $r_{t}$ equal to the splash-back radius of the halo (Diemer & Kravtsov, 2014; Adhikari et al., 2014; More et al., 2015), and for subhalos we choose $r_{t}$ based on the subhalo mass and three-dimensional position inside the host (see Gilman et al. (2020)).
The central density $\rho_{s}$ determines the lensing efficiency of a halo of a fixed total mass. The delayed onset of structure formation in WDM models suppresses the central density of halos with mass below $m_{\rm{hm}}$ (Bose et al., 2016; Ludlow et al., 2016). The concentration-mass relation establishes the connection between $\rho_{s}$ and halo mass through the concentration parameter, $c$. We use the concentration-mass relation presented by Diemer & Joyce (2019), and implemented with the software colossus (Diemer, 2018), to assign concentrations to CDM halos, and compute the concentrations in WDM according to (Bose et al., 2016)
@@
We generate line-of-sight halos (subhalos) with (infall) masses in the range $10^{6}-10^{10}M_{\odot}$. Halos more massive than $10^{10}M_{\odot}$ would likely contain a luminous galaxy, in which case we would insert these objects explicitly into the lens model. Halos less massive than $10^{6}M_{\odot}$ lie below the sensitivity threshold of our data given our assumptions regarding the background source size (see the next sub-section).
In the dark matter inference, we sample $\Sigma_{\rm{sub}}$ from a log-uniform prior $\log_{10}\Sigma_{\rm{sub}}\in\mathcal{U}\left(-2.5,-1.0\right)$. This is a broad and uninformative prior that we use to reveal any degeneracies between $\Sigma_{\rm{sub}}$ and $m_{\rm{hm}}$, and to understand to what degree imaging data can break these degeneracies. We sample $\log_{10}m_{\rm{hm}}\in\mathcal{U}\left(4.0,10.0\right)$. For both the halo mass function and concentration-mass relation, values of $m_{\rm{hm}}<10^{5}M_{\odot}$ result in mass functions and concentration-mass relations that deviate the CDM prediction below the estimated sensitivity threshold of our data; thus, we can consider these realizations as consistent with CDM. We fix the values of all other parameters introduced in this section to the stated values.
@@ 4.2.1 Creating mock lens mass profiles
To test the methodology discussed in Section 2 we create a sample of 25 mock lens systems with properties broadly comparable to the know population of such systems (Auger et al., 2010; Oguri & Marshall, 2010) with which to test the lens modeling approach outlined in Section 2. The mocks have lens (source) redshifts in the range 0.3 - 0.9 (0.9 - 3.0). We model the main deflector galaxy as a elliptical power-law (EPL) mass profile with an Einstein radius set to $1\ \rm{arcsec}$ for each system, axis ratios in the range 0.50 - 0.95, and logarithmic mass profile slopes $\gamma$ drawn from a Gaussian prior $\mathcal{N}\left(2.0,0.1\right)$. We apply external shear across the main lens plane with a random orientation and a strength $\gamma_{\rm{ext}}$ in the range 0.02 - 0.16.
@@
We base our implementation of the multipole perturbations on the observed properties of 847 elliptical galaxies presented by Hao et al. (2006). The shape of the iso-density contours inferred from the light (and assuming light traces mass) can be related to the physical amplitude of the perturbation by $a_{\rm{m(phys)}}=a_{\rm{m}}\times\frac{\theta_{E}}{\sqrt{q}}$, where $\theta_{\rm{E}}$ is the Einstein radius and $q$ is the axis ratio of the ellipse. The third ($m=3$) and fourth-order ($m=4$) moments make the dominant contribution to the multipole expansion of galaxy shapes (Bender et al., 1989). Hao et al. (2006) find $a_{\rm{3}}$ distributed as $\mathcal{G}\left(0.0,0.005\right)$ with values of $\phi_{3}$ uncorrelated with the position angle of the underlying mass profile. Hao et al. (2006) also measure $a_{4}\in\mathcal{G}\left(0.0,0.01\right)$ with an orientation that tends to align with the orientation of the underlying mass profile. When the $a_{4}$ orientation aligns with the orientation of the underlying ellise, the profile appears boxy ($a_{4}<0$) or disky ($a_{4}>0$). The observed amplitudes $a_{3}$ and $a_{4}$ have no apparent correlation.
Based on the findings by Hao et al. (2006), we include the dominant $m=3$ and $m=4$ terms and choose priors for their observed amplitudes $a_{3}\in\mathcal{G}\left(0.0,0.005\right)$ and $a_{4}\in\mathcal{G}\left(0.0,0.01\right)$. We enforce alignment between the $a_{4}$ orientation and the underlying EPL profile (producing boxy or disky contours), and sample $\phi_{3}\in\mathcal{U}\left(-\pi/6,\ \pi/6\right)$.
@@ 4.2.2 Modeling of the lens mass profile
When modeling the mock lenses, all of the parameters that describe the EPL profile are left free to vary while reconstructing the imaging data, as well as the strength and position angle of the external shear. For each lens, we draw samples of $a_{3}$, $a_{4}$, and $\phi_{3}$ from the same priors as those used to create the simulated datasets. We keep the amplitudes of the multipole terms fixed to the values drawn from their respective priors while reconstructing the imaging data, and eventually constrain $a_{3}$ and $a_{4}$ when down-selecting on the imaging data summary statistic, as discussed in Section 2.2. We have experimented with allowing $a_{3}$ and $a_{4}$ to vary freely while reconstructing the imaging data, but find that this causes their inferred amplitudes to take on unphysical amplitudes, well beyond the range of a reasonable expectation for their values. This unexpected behavior could arise from degeneracies in the imaging data likelihood between multipole perturbations to the lens model and full populations of dark matter halos, as well as limitations associated with the PSF and source reconstruction. During the lens modeling, we include a Gaussian prior on $\gamma$, the logarithmic slope of the main deflector mass profile, with a mean of -2.0 and standard deviation of 0.2.
@@ 4.3.1 Creating mock lens and source light profiles
We parameterize the main deflector light as a circular Sersic profile. The source light model includes two components: the lensed quasar and its host galaxy. For the quasar, we assume flux ratios measured from the warm dust region now observable with JWST (Nierenberg et al., 2023) with a physical size $\sim 1-10\ \rm{pc}$. To create the mocks, we model the quasar emission as a circular Gaussian with a full-width at half maximum sampled from a uniform prior $\mathcal{U}\left(1-10\right)\rm{pc}$. We assume a flux ratio measurement precision of $3\%$, and an astrometric precision in the relative quasar image positions of 5 milli-arcseconds, based on the recent measurements (Nierenberg et al., 2023).
To create realistic lensed arcs, we place the spiral galaxy shown in Figure 4 at the source redshift for each system. We extract this source from the COSMOS survey catalog (Koekemoer et al., 2007) using the software package paltas (Wagner-Carena et al., 2023). We place this same source behind each of the 25 mock deflectors to simulate the imaging data. This particular galaxy was selected because it exhibits enough morphological complexity (spiral arms) to warrant a small-scale basis set expansion of its light profile to reproduce a lensed image. Many real strong lens systems exhibit this level of complexity in their inferred source light profiles, and we include this feature in our simulations to determine whether our methodology yields unbiased inferences of substructure properties with a complex source morphology.
@@ 4.3.2 Modeling of lens and source light profiles
We model the main deflector light as an elliptical Sersic profile. During the inference we parameterize the quasar emission region as a circular Gaussian, and sample its size uniformly $\mathcal{U}\left(1,10\right)\ \rm{pc}$ for each system. We model the source with an ellpitical Sersic profile with additional small-scale complexity implemented through shapelet basis functions (Birrer et al., 2015). An integer $n_{\rm{max}}$ determines the degree of complexity these functions can describe, with the complexity of the reconstructed source image with increasing $n_{\rm{max}}$.
@@
where $k$ is the number of model parameters, $n$ represents the number of data points, and $\max\left[\mathcal{L}\right]$ represents the maximum likelihood of the data given the model. The BIC statistic compensates between obtaining a better fit to data and over-fitting a model. For each mock system, we found the BIC reaches a minimum for a source model consisting of an elliptical Sersic profile plus shapelets at $n_{\rm{max}}=10$. For accepted realizations, we obtain a reduced $\chi^{2}$ per degree of freedom $\chi^{2}/\rm{DOF}\approx 1.2$. Appendix A provides additional discussion regarding the reconstruction of the source light in our simulations.
@@ Acknowledgments
Computations performed for this analysis were performed on the Midway2 and Midway3 clusters hosted by the University of Chicago, the Niagara supercomputer at the SciNet HPC Consortium, and the Hoffman2 Shared Cluster provided by the UCLA Institute for Digital Research and Education’s Research Technology Group. SciNet is funded by the Canada Foundation for Innovation; the Government of Ontario; Ontario Research Fund - Research Excellence; and the University of Toronto.

### 10 Mar 2024 v2

@@ 1 Introduction
In search of new physics, observational probes of dark matter structure from dwarf galaxies (Kim et al., 2018; Correa, 2021; Nadler et al., 2021; Bechtol et al., 2022; Dekker et al., 2022; Akita & Ando, 2023; Slone et al., 2023; Nadler et al., 2024), stellar streams (Bovy et al., 2017; Banik et al., 2018; Bonaca et al., 2019; Banik et al., 2021), and gravitational lensing (Dalal & Kochanek, 2002; Vegetti et al., 2014; Nierenberg et al., 2014; Hezaveh et al., 2016; Birrer et al., 2017; Nierenberg et al., 2017; Vegetti et al., 2018; Gilman et al., 2020, 2021; He et al., 2022; Sengül et al., 2022; Wagner-Carena et al., 2023; Dike et al., 2023; Powell et al., 2023; Dhanasingham et al., 2023; Keeley et al., 2023; Gilman et al., 2023; Mondino et al., 2023; Nightingale et al., 2024) characterize the properties of substructure on sub-galactic scales. If the dark matter has no coupling to the standard model besides gravity, these cosmic probes of dark matter constitute the only experiments with which to investigate its properties (see the reviews by Drlica-Wagner et al., 2022; Bechtol et al., 2022).
@@
We perform lensing calculations using the open-source software package lenstronomy [^1] https://github.com/lenstronomy/lenstronomy (Birrer & Amara, 2018; Birrer et al., 2021). We generate the populations of dark matter substructure using the open-source software pyHalo [^2] https://github.com/dangilman/pyHalo (Gilman et al., 2020). To perform the forward modeling calculations that interface between lenstronomy and pyHalo, we use the new open-source software samana [^3] https://github.com/dangilman/samana. We assume a flat cosmology with $\Omega_{\rm{m}}=0.32$, $\sigma_{8}=0.81$ and $H_{0}=67.4\ \rm{km}\ \rm{s^{-1}}\ \rm{Mpc^{-1}}$ (Planck Collaboration et al., 2020).
@@ 4.1 Modeling of dark matter subhalos and line-of-sight halos
where $\rho_{s}$ is a characteristic central density, $r_{s}$ is the scale radius, and $r_{t}$ implements a tidal radius. For field halos, we set $r_{t}$ equal to $r_{\rm{50}}$, which is comparalbe to the splash-back radius of a halo (Diemer & Kravtsov, 2014; Adhikari et al., 2014; More et al., 2015). We truncated subhalo density profiles based on the mass and three-dimensional position inside the host (see Gilman et al. (2020)).
The central density $\rho_{s}$ determines the lensing efficiency of a halo of a fixed total mass. The delayed onset of structure formation in WDM models suppresses the central density of halos with mass below $m_{\rm{hm}}$ (Bose et al., 2016; Ludlow et al., 2016). The concentration-mass relation establishes the connection between $\rho_{s}$ and halo mass through the concentration parameter, $c$. We use the concentration-mass relation presented by Diemer & Joyce (2019), and implemented with the software colossus (Diemer, 2018), to assign concentrations to CDM halos. We compute the concentrations in WDM according to (Bose et al., 2016)
@@
We generate line-of-sight halos (subhalos) with (infall) masses in the range $10^{6}-10^{10}M_{\odot}$. Halos more massive than $10^{10}M_{\odot}$ would likely contain a luminous galaxy, in which case we would include these objects in the lens macromodel. Halos less massive than $10^{6}M_{\odot}$ lie below the sensitivity threshold of our data given our assumptions regarding the background source size (see the next sub-section).
In the dark matter inference, we sample $\Sigma_{\rm{sub}}$ from a log-uniform prior $\log_{10}\Sigma_{\rm{sub}}\in\mathcal{U}\left(-2.5,-1.0\right)$. This is a broad and uninformative prior that we use to reveal any degeneracies between $\Sigma_{\rm{sub}}$ and $m_{\rm{hm}}$, and to understand to what degree imaging data can break these degeneracies. We sample $\log_{10}m_{\rm{hm}}\in\mathcal{U}\left(4.0,10.0\right)$. For both the halo mass function and concentration-mass relation, values of $m_{\rm{hm}}<10^{5}M_{\odot}$ result in mass functions and concentration-mass relations that deviate from the CDM prediction below the estimated sensitivity threshold of our data; thus, we can consider these realizations as consistent with CDM. We fix the values of all other parameters introduced in this section to the stated values.
@@ 4.2.1 Creating mock lens mass profiles
To test the methodology discussed in Section 2 we create a sample of 25 mock lens systems with properties broadly comparable to the know population of such systems (Auger et al., 2010; Oguri & Marshall, 2010). The mocks have lens (source) redshifts in the range 0.3 - 0.9 (0.9 - 3.0). We model the main deflector galaxy as a elliptical power-law (EPL) mass profile with an Einstein radius set to $1\ \rm{arcsec}$ for each system, axis ratios in the range 0.50 - 0.95, and logarithmic mass profile slopes $\gamma$ drawn from a Gaussian prior $\mathcal{N}\left(2.0,0.1\right)$. We apply external shear across the main lens plane with a random orientation and a strength $\gamma_{\rm{ext}}$ in the range 0.02 - 0.16.
@@
We base our implementation of the multipole perturbations on the observed properties of 847 elliptical galaxies presented by Hao et al. (2006). The shape of the iso-density contours inferred from the light (and assuming light traces mass) can be related to the physical amplitude of the perturbation by $a_{\rm{m(phys)}}=a_{\rm{m}}\times\frac{\theta_{E}}{\sqrt{q}}$, where $\theta_{\rm{E}}$ is the Einstein radius, $q$ is the axis ratio of the ellipse and $a_{m}$ represents the observable deviation from ellipticity of the light. The third ($m=3$) and fourth-order ($m=4$) moments make the dominant contribution to the multipole expansion of galaxy shapes (Bender et al., 1989). Hao et al. (2006) find $a_{\rm{3}}$ distributed as $\mathcal{G}\left(0.0,0.005\right)$ with values of $\phi_{3}$ uncorrelated with the position angle of the underlying mass profile. Hao et al. (2006) also measure $a_{4}\in\mathcal{G}\left(0.0,0.01\right)$ with an orientation that tends to align with the orientation of the underlying mass profile. When the $a_{4}$ orientation aligns with the orientation of the underlying ellipse the profile appears boxy ($a_{4}<0$) or disky ($a_{4}>0$). The observed amplitudes $a_{3}$ and $a_{4}$ have no apparent correlation.
Based on the findings by Hao et al. (2006), when creating mock deflectors we include the dominant $m=3$ and $m=4$ terms and use priors for their observed amplitudes $a_{3}\in\mathcal{G}\left(0.0,0.005\right)$ and $a_{4}\in\mathcal{G}\left(0.0,0.01\right)$. Based on the measurements by Hao et al. (2006) and Oh et al. (in prep), we enforce alignment between the $a_{4}$ orientation and the underlying EPL profile (producing boxy or disky contours), and sample $\phi_{3}\in\mathcal{U}\left(-\pi/6,\ \pi/6\right)$.
@@ 4.2.2 Modeling of the lens mass profile
When modeling the mock lenses, all of the parameters that describe the EPL profile are left free to vary while reconstructing the imaging data, as well as the strength and position angle of the external shear. We draw $a_{3}$, $a_{4}$, and $\phi_{3}$ from the same priors as those used to create the simulated datasets. During the reconstruction of the imaging data for each realization, we keep the multipole terms fixed to the values drawn from their respective priors. We eventually constrain these parameters when down-selecting on the imaging data and flux ratio summary statistics, as discussed in Section 2.2.
We have experimented with allowing $a_{3}$ and $a_{4}$ to vary freely while reconstructing the imaging data, but find that this causes their inferred amplitudes to take on unphysical values. This unexpected behavior could arise from degeneracies in the imaging data likelihood between multipole perturbations to the lens model and full populations of dark matter halos, as well as limitations associated with the PSF and source reconstruction. During the lens modeling, we include a Gaussian prior on $\gamma$, the logarithmic slope of the main deflector mass profile, with a mean of -2.0 and standard deviation of 0.2.
@@ 4.3.1 Creating mock lens and source light profiles
We parameterize the main deflector light as a circular Sérsic profile (Sérsic, 1963). The source light model includes two components: the lensed quasar and its host galaxy. For the quasar, we assume flux ratios measured from the warm dust region now observable with JWST (Nierenberg et al., 2023) with a physical size $\sim 1-10\ \rm{pc}$. To create the mocks, we model the quasar emission as a circular Gaussian with a full-width at half maximum sampled from a uniform prior $\mathcal{U}\left(1-10\right)\rm{pc}$. We assume a flux ratio measurement precision of $3\%$, and an astrometric precision in the relative quasar image positions of 5 milli-arcseconds, based on the recent measurements (Nierenberg et al., 2023).
To create realistic lensed arcs, we place the spiral galaxy shown in Figure 4 at the source redshift for each of the 25 mock deflectors. We extract this source from the COSMOS survey catalog (Koekemoer et al., 2007) using the software package paltas (Wagner-Carena et al., 2023). This particular galaxy was selected because it exhibits enough morphological complexity (spiral arms) to warrant a small-scale basis set expansion of its light profile to reproduce a lensed image. Many real strong lens systems exhibit this level of complexity in their inferred source light profiles, and we include this feature in our simulations to determine whether our methodology yields unbiased inferences of substructure properties with a complex source morphology.
@@ 4.3.2 Modeling of lens and source light profiles
We model the main deflector light as an elliptical Sérsic profile. We model the quasar emission region as a circular Gaussian, and sample its size uniformly $\mathcal{U}\left(1,10\right)\ \rm{pc}$ for each system. We model the quasar host galaxy with an ellpitical Sérsic profile with additional small-scale complexity implemented through shapelet basis functions (Birrer et al., 2015). An integer $n_{\rm{max}}$ determines the degree of complexity these functions can describe, with the complexity of the reconstructed image with increasing $n_{\rm{max}}$.
@@
where $k$ is the number of model parameters, $n$ represents the number of data points, and $\max\left[\mathcal{L}\right]$ represents the maximum likelihood of the data given the model. The BIC statistic compensates between obtaining a better fit to data and over-fitting a model. To calculate the BIC, we fit a smooth lens model (i.e. a lens model without substructure) to each mock. For each mock system, we found the BIC reaches a minimum for a source model consisting of an elliptical Sérsic profile plus shapelets at $n_{\rm{max}}=10$. With this source model we obtain a reduced $\chi^{2}$ per degree of freedom $\chi^{2}/\rm{DOF}\approx 1.2$ for the lens models we accept based on the imaging data likelihood. Appendix A provides additional discussion regarding the reconstruction of the source light in our simulations.
@@ Acknowledgments
Computations performed for this analysis were performed on the Midway2 and Midway3 clusters hosted by the University of Chicago’s Research Computing Center, the Niagara supercomputer at the SciNet HPC Consortium, and the Hoffman2 Shared Cluster provided by the UCLA Institute for Digital Research and Education’s Research Technology Group. SciNet is funded by the Canada Foundation for Innovation; the Government of Ontario; Ontario Research Fund - Research Excellence; and the University of Toronto.